OpenStudy (jemson):

An employer pays $26 for worked days and reduces $7 for being absent. If an employee received $829 after 56 days, then how many days was he absent?

OpenStudy (nincompoop):

s = 26d - 19a 829 = 26(56) + 19a 829 - 1456 = 19a -33 = a where s: salary d: days work a: absent; it says "reduces $7 for being absent" that may mean that instead of 26, the employee receives 19 only. Paid day off LAUGHING OUT LOUD the employee can be absent for the entire year and not work and still receive a lot of money.

OpenStudy (jemson):

but why -19a??? it should be -7a, as he receives $7 extra for all the days.

OpenStudy (jemson):

33*19=627(Absent day salary) & 23*26=598(Present days salary) . Total 627+598 = 1225, which is more than what he gets in 56 days..